# TrustGeoGen: Scalable and Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

### v1

@@ 5 Conclusion and Further Discussion
Conclusion. We present TrustGeoGen, a multimodal-integrated and scalable geometric data engine that automatically generates fully verified geometric reasoning data. The resulting GeoTrust dataset demonstrates validated effectiveness and generalization capabilities, with its trustworthy reasoning data showing significant advantages over conventional pseudo-labels. Notably, the GeoTrust-test exposes critical limitations of existing MLLMs in handling complex geometric reasoning tasks. GenGen represents a crucial first step towards trustworthy geometric problem solving, establishing a new paradigm for generating rigorous data.
@@
Trustworthy Geometric Problem Solving. Models should produce verifiable reasoning steps, not merely correct answers. Achieving this demands both reliable data and robust evaluation mechanisms are required. Leveraging the generated data from TrustGeoGen, further study can focus on autoformalization approachs that translate natural language reasoning into formalized steps, enabling automated verification of each deduction’s logical validity. Moreover, TrustGeoGen can dynamically generate evaluation data to prevent prior data leakage during evaluation, ensuring higher reliability in unseen geometric scenarios.
@@
Formal Enhanced Reasoning. TrustGeoGen’s formalized reasoning environment enables the generation of trustworthy geometric data by constructing rigorous reasoning graphs that ensure logical correctness at each step. These graphs provide a structured foundation for exploring diverse mathematical reasoning strategies through tailored sampling methods: the current work implements multi-solution and self-reflection traceback data, while future extensions could incorporate the idea of reverse thinking and categorical discussion, etc. These ideas could be further enhanced by integrating alternative training approaches (e.g. RL) to generalize to other reasoning scenarios.
Appendix
Appendix A Related Work
Advances in Multimodal Large Language Model.
Recent advances in Large Language Models (LLMs) have demonstrated remarkable progress in linguistic intelligence, achieving human-level capabilities in various applications [26, 33, 34, 32]. Building upon this foundation, the research community has focused on extending these text-based architectures to process visual information, leading to the emergence of sophisticated Multimodal Large Language Models (MLLMs) [4, 1, 29]. These integrated systems typically employ modality alignment mechanisms, such as Q-former [17] and linear layers [19], to establish connections between visual representations and text embeddings. Although current MLLMs show strong performance in standard vision-language benchmarks [13, 39, 38, 22, 16], their effectiveness decreases significantly when processing mathematical visualizations that require reasoning. Recent specialized approaches address this limitation through targeted training strategies. For example, geometric reasoning capabilities have been enhanced through domain-specific datasets containing annotated diagrams [14, 46].
Model Name Model / API Version
GPT-4o [24] gpt-4o-2024-11-20
Claude-3-7-Sonnet [3] claude-3-7-sonnet-20250219
DeepSeek-R1 [10] deepseek-r1
OpenAI-o1 [25] o1-2024-12-17
Table 4: Model / API versions used for evaluation across different MLLMs.
Geometric Problem Solving.
As a challenging task, geometric problem solving requires understanding diagrams, interpreting symbols, and performing complex reasoning. While MLLMs have shown remarkable proficiency in human-level reasoning and generation capabilities, they struggle with automatic geometric problem solving due to their pre-training on natural images and texts, as well as the lack of automated verification in the problem-solving process. To address this limitation, researchers have proposed various approaches to improve the understanding of geometric images and corpora, including unimodal pre-training, vision-language alignment, and visual instruction tuning [40, 46, 14, 6, 7, 44, 15, 16]. Another major approach involves the use of formalized solvers to tackle geometric problems. While these solvers [35, 9, 31] are capable of addressing challenges at the level of the International Mathematical Olympiad (IMO), they require precise modeling of geometric problems to fit the solver’s language. This necessity introduces significant obstacles in terms of practical application and generalization, limiting the usability in real-world scenarios.
Appendix B Versions of Closed-source Model
During the evaluation on GeoTrust-test, we utilize some closed-source commercial models, with the specific version detailed in Table 4.
Appendix C More Visualization Example
In Fig. 5, we provide one more visualization example in GeoTrust with 94 reasoning steps and 0.95 premises ratio.
Figure 5: One visualization example in GeoTrust.

### v2

@@ 6 Conclusion and Further Discussion
Conclusion. We have introduced TrustGeoGen, a multi-modal, integrated, and formal-verified engine for automatically generating fully geometric reasoning data. The resultant GeoTrust dataset demonstrates proven effectiveness and generalization capabilities. Furthermore, our benchmark, GeoTrust-test, reveals critical limitations in existing multimodal large language models (MLLMs) when handling complex geometric reasoning tasks. Notably, the proposed "connection thinking" bridges the gap between formal and natural reasoning, offering a valuable blueprint for synthesizing model training data with formal engines. In parallel, the "thinking template" mechanism enriches the diversity of the synthetic data, providing greater information gain and further enhancing the models’ reasoning abilities. TrustGeoGen represents a crucial first step toward trustworthy geometric problem-solving, establishing a new paradigm for generating rigorous and reliable data.
Figure 9: Ablation Study of Thinking Template on Geometry3K
Further Discussion. We hope that future research will develop from the following two perspectives:
@@
Trustworthy Geometric Problem Solving: Models should produce verifiable reasoning steps, not merely correct answers. Achieving this demands both reliable data and robust evaluation mechanisms are required. Leveraging the generated data from TrustGeoGen, further study can focus on autoformalization approachs that translate natural language reasoning into formalized steps, enabling automated verification of each deduction’s logical validity. Moreover, TrustGeoGen can dynamically generate evaluation data to prevent prior data leakage during evaluation, ensuring higher reliability in unseen geometric scenarios.
@@
Formal Enhanced Reasoning: TrustGeoGen’s formalized reasoning environment enables the generation of trustworthy geometric data by constructing rigorous reasoning graphs that ensure logical correctness at each step. These graphs provide a structured foundation for exploring diverse mathematical reasoning strategies through tailored sampling methods: the current work implements multi-solution and self-reflection traceback data, while future extensions could incorporate the idea of reverse thinking and categorical discussion, etc. These ideas could be further enhanced by integrating alternative training approaches (e.g. RL) to generalize to other reasoning scenarios.
Acknowledgements
The research was supported by Shanghai Artificial Intelligence Laboratory, a locally commissioned task from the Shanghai Municipal Government, National Natural Science Foundation of China (Grant No. 92370201 and 62222607), and the Shanghai Municipal Science and Technology Major Project (Grant No. 22DZ1100102).